OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 · x + 7 · 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (anonymous):

\[a \left( b+c \right)=a*b+a*c\]

OpenStudy (anonymous):

@AccessDenied @math_genius123

OpenStudy (anonymous):

@surjithayer

OpenStudy (linda3):

Okay- so do you know how to get the answer for 7(5+9) ?

OpenStudy (anonymous):

not really i try to practice on it.

OpenStudy (linda3):

okay thats fine I'll lead you through it

OpenStudy (anonymous):

no wait i do

OpenStudy (anonymous):

7(5+9)= 7(14)

OpenStudy (anonymous):

98

OpenStudy (linda3):

(7x5) + (7x9) = 35 + 63 =98 good job!

OpenStudy (anonymous):

thanks whats the next step

OpenStudy (linda3):

So now what is 7 · x + 7 · 9 ? it would be 7x + 63

OpenStudy (linda3):

so the problem looks like this 98 = 7x +63

OpenStudy (anonymous):

okay

OpenStudy (linda3):

Now your goal is to get x alone on one side. In order to do this you have to subtract 63 by both sides and then you'll get 35 = 7x now you divide 7 by both sides and you'll get 35 divide by 7 is?

OpenStudy (anonymous):

5

OpenStudy (linda3):

Good job!!! :D

OpenStudy (anonymous):

so the answer is 5?

OpenStudy (linda3):

yes you are correct

OpenStudy (anonymous):

Thanks!

OpenStudy (linda3):

Your most very welcome!